Based on the reaction, Sn + Cu2+ ⇌ Sn2+ + Cu+ (in acidic medium)

Calculate the following values:

a.      E°cell

b.     ΔG°

c.      K

d.     E (when [Cu2+] = 0.0200 M, [Sn2+] = 0.0100 M, and [Cu+] = 0.0300 M)

 


1
Expert's answer
2022-01-17T11:47:02-0500

Question:

Based on the reaction, Sn + Cu2+ ⇌ Sn2+ + Cu(in acidic medium)

Calculate the following values:

a.      E°cell

b.     ΔG°

c.      K

d.     E (when [Cu2+] = 0.0200 M, [Sn2+] = 0.0100 M, and [Cu+] = 0.0300 M)


Solution:

The balanced reaction is:

Sn + 2Cu2+ ⇌ Sn2+ + 2Cu+

The half reactions occurring on the electrodes are:

Sn - 2e- "\rightarrow" Sn2+

Cu2+ + e- "\rightarrow" Cu+

Thus, the first reaction is the oxidation ( so the electrode is the anode) and the second reaction is the reduction (so the electrode is cathode).

The standard cell potential E°cell is the difference of the standard reduction potential of the electrodes:

"E\u00b0_{cell} = E\u00b0_{red, cathode} - E\u00b0_{red, anode}"

The standard potentials for the half reactions shown above are:

Sn - 2e- "\rightarrow" Sn2+ "E\u00b0_{red, anode} = -0.137" V

Cu2+ + e- "\rightarrow" Cu+ "E\u00b0_{red, cathode} = 0.340" V

Therefore, the standard potential of the cell is:

"E\u00b0_{cell} = 0.340 - (-0.137) = 0.477" V


The change in free energy "\Delta G" is related to the standard potential as follows:

"\u2206G = -nFE\u00b0"

where "nF" is the charge transferred during the reaction, 2·96485 J/V. Therefore:

"\u2206G = -2\u00b796485\u00b70.477 = -92" kJ.


The relationship between the change in free energy and the equilibrium constant is:

"\u2206G=-RTlnK"

"K = \text{exp}(-\frac{∆G}{RT}) = \text{exp}(\frac{92047}{8.314·298}) = 1.4·10^{16}"


According to the Nernst equation, the electrochemical cell potential depends on the concentrations of the reactants and products:

"E = E° - \frac{RT}{nF}\text{log}\frac{[Sn^{2+}][Cu^+]^2}{[Cu^{2+}]^2}"

"E = 0.477 - \frac{8.314·298}{2·96485}\text{log}\frac{0.01·0.03^2}{0.02^2} = 0.498" V.


Answer:

a. E°cell = 0.477 V

b.     ΔG° = -92 kJ

c.     K = 1.4·1016

d.     E (when [Cu2+] = 0.0200 M, [Sn2+] = 0.0100 M, and [Cu+] = 0.0300 M) = 0.498 V